# FootFormer: Estimating Stability from Visual Input

### v1

@@ 5.4 Foot Contact
Estimating foot contact (FC), or whether a foot or specific parts of the foot are in contact with the ground plane, is essential for applications in locomotion analysis, rehabilitation, graphics, and animation. We compare FootFormer against FPP-Net [22] and UP [17] (Table 1) on their respective datasets to assess effectiveness in contact prediction. Table 3 presents the precision (prec.), recall, F1-score, and IoU evaluation scores.
@@
UP[17] UnderPressure $0.936^{\dagger}$ $^{\dagger} 0.9450.945^{\dagger}$ $0.896^{\dagger}$
@@
Table 3: Foot contact estimation results. We train and evaluate on the MMVP [22] and UnderPressure [17] datasets and compare against their respective baseline models. Bold indicates the best (highest metric); denotes a statistically significant difference from Ours under paired t-test (p < 0.05).
On the MMVP dataset, FootFormer achieves significantly better precision and equivalent F1, IoU, and recall. Moreover, our model achieves statistically significant improvements over UP across all 4 metrics. Exact p-values are provided in the Supplementary Material. Our results for FPP-Net on the MMVP dataset may be superior to those reported in the original paper [22] as we fully retrain and evaluate with OpenPose [1] keypoints instead of their original keypose extraction method [54].
@@ 5.5 Stability Components
Our goal is comprehensive stability analysis. To this end, we evaluate estimates of three stability components, CoP, CoM and BoS, that form the foundation for quantifying human postural stability and balance [56, 57, 42]. Figure 3 depicts foot pressure with these stability components and two stability components used in their calculation.
@@
Center of Mass (CoM): CoM represents the weighted average position of body mass and is a crucial factor in a person’s ability to maintain balance. Unlike CoP and BoS, CoM is directly regressed from keypoint sequences. We compare against two baseline methods, CoMNet [20], a fully connected network (Table 1), and Dempster’s method [18, 58, 59], a classical anthropomorphic method that estimates CoM from weighted sums of segmental centers of mass across the body. For CoM evaluation, Figure 4(a,b) presents both mean and median L2 errors reported in millimeters (mm). FootFormer demonstrates statistically significant improvements over both classical (Dempster’s method [18, 58, 59]) and learned (CoMNet [20]) baselines.
@@
Figure 4: Performance comparison on  PSU-TMM100 [19]. (a,b) Mean and median CoM L2 error across subjects. (c) CoP L2 error and (d) BoS IoU across varying pressure thresholds. FootFormer achieves statistically significantly better results across all stability components (see Supplementary Material for exact p-values).
Center of Pressure (CoP): CoP is calculated as the weighted mean of the pressure elements in the XY ground plane. Accurate estimation of CoP is essential for understanding balance, as shifts in CoP can indicate changes in stability or an impending need for postural adjustment [56]. Base of Support (BoS): BoS represents the area under the feet that supports the body. Estimating BoS is critical for determining the boundaries within which an individual can maintain balance [57]. We follow [20] and perform foot localization to align the predicted and GT pressure maps with the floor plane using 3D-triangulated keypoints from two viewpoints. Foot position is estimated by foreshortening the pressure map based on dorsiflexion and plantar flexion angles, and rotating it according to ankle orientation ensuring spatial consistency.
@@ 5.6 Stability
Moving beyond simple kinematic estimates, the multiple outputs of FootFormer enable us to directly estimate stability. We calculate two popular measures of postural stability (Figure 3). First, we estimate CoM-CoP defined as $\|CoM-CoP\|_{2}$ or the Euclidean distance from the 2D CoM projected onto the floor plane to CoP [56]. Typically, the further apart these two points are, the greater the potential for becoming unstable. Second, we measure $\|CoM-BoS_{nearest}\|_{2}$ or the Euclidean distance from the 2D CoM to the boundary of the BoS [57]. Intuitively, the CoM-BoS captures the magnitude of instability.
@@
CoM-CoP CoMNet+PNS [20] $46.00\pm 22.0^{\dagger}$ $29.86\pm 13.2^{\dagger}$
@@
Table 4: Stability quantification results on PSU-TMM100 [19]. CoM-CoP and CoM-BoS are reported in mm. Bold indicates the best (lowest error); denotes a statistically significant difference from Ours under paired t-test (p < 0.05).
We compare with Scott et al. [20] who use CoMNet to estimate CoM and PNS [19] to regress foot pressure (Table 1). Table 4 follows [20] and reports the mean $\pm$ std and median $\pm$ rSTD for CoM-CoP and CoM-BoS error in mm, where rSTD represents robust standard deviation calculated as the median absolute deviation from the median, multiplied by 1.4826 [60]. Error is computed as the absolute distance between predicted and ground-truth positions derived from the insole sensors and MoCap system used in PSU-TMM100. FootFormer achieves statistically significantly improvements in both stability metrics compared to the combined multi-model approach of CoMNet+PNS. We believe this validates the efficacy of learning coupled motion dynamics within a unified structure, as our joint optimization approach outperforms separate models trained independently for each component.
@@ 5.7 Ablation Experiment
Pose Embedding: We compare our learnable GCN against a 1D CNN [22, 17] and a linear MLP [19]. Temporal Modeling: We replace the STT with a standard transformer and GRU [22] to assess the efficacy of the spatiotemporal attention mechanism. Contact Conditioning: We evaluate the pressure decoder with and without the contact-based gating to investigate the cross-modal alignment benefits.
@@ A.1 Detailed Architecture Specifications
To adapt the UP [17] model to PSU-TMM100, we make simple changes to resize the input and final regression layer to fit PSU-TMM100’s pose input and insole pressure maps, respectively; all other network components are maintained as is. Similarly, we modify FPP-Net’s [22] first layer to fit the joint scheme present in the data. We then adapt the network’s pressure and contact regressor to fit the insole shape and contact regions (like that of FootFormer).
@@ Appendix B Additional Foot Pressure Estimation Results
Figure 6: KLD foot pressure estimation results on PSU-TMM100 [19] across each subject with mean and median values. We train and evaluate on both 2D and 3D detected keypoints. Lower is better. Statistical significance values are presented in Table 5.
We report the full comparison with Scott et al. [19], UP [17], and FPP-Net [22] on the PSU-TMM100 for each subject in Figure 6. Table 5 reports statistical significance values of paired t-tests comparing FootFormer with the baseline models.
@@
PNS [19] 2.82 $\pm$ 0.86($1.45\mathrm{e}{-04}$) 2.68 $\pm$ 0.94($9.38\mathrm{e}{-03}$)
FPP-Net [22] 1.40 $\pm$ 0.32   ($2.93\mathrm{e}{-01}$) 1.60 $\pm$ 0.48($1.61\mathrm{e}{-02}$)
UP [17] 1.45 $\pm$ 0.35   ($6.06\mathrm{e}{-02}$) 1.50 $\pm$ 0.35($1.59\mathrm{e}{-02}$)
@@
Table 5: Comparison of FootFormer (Ours) with baselines on PSU-TMM100 [19]. Each entry reports mean $\pm$ std KLD and the paired t-test p-value vs. Ours. Bold indicates the best (lowest KLD); denotes a statistically significant difference from Ours (p < 0.05). Per-subject results are shown in Fig. 6.
@@ 11 Full-Body Stretches 2216
  • Foot pressure maps: Recorded using Tekscan F-Scan 7.0 [48] insole sensors at 50 Hz. Each foot has a high-resolution prexel grid of size $60\times 21$, capturing pressure intensities in kilopascals.
  • 2D and 3D body pose: 2D joints are extracted using OpenPose [1] BODY25, while 3D joints are reconstructed via stereo triangulation from the two camera views (like in PSU-TMM100).
@@
Figure 7: Qualitative comparison of predicted foot pressure maps across five OM activities for FootFormer (Ours), PNS [19], FPP-Net [22], and UP [17]. Column headers indicate action; row headers indicate source view or prediction method.
@@ C.1 Additional Ordinary Movements Evaluation
PNS [19] 3.53 $\pm$ 1.23 ($1.65\mathrm{e}{-04}$) 2.59 $\pm$ 0.94 ($4.35\mathrm{e}{-03}$)
FPP-Net [22] 1.52 $\pm$ 0.37 ($6.89\mathrm{e}{-01}$) 1.54 $\pm$ 0.34 ($9.04\mathrm{e}{-01}$)
UP [17] 1.69 $\pm$ 0.34 ($1.28\mathrm{e}{-01}$) 1.48 $\pm$ 0.41 ($6.66\mathrm{e}{-01}$)
@@ Appendix D Additional Foot Contact Estimation Results
Table 8 reports statistical significance values of paired t-tests comparing FootFormer with FPP-Net and UP on the MMVP [22] and UnderPressure [17] datasets respectively.
@@
FPP-Net [22] MMVP $0.635^{\dagger}$($1.28\mathrm{e}{-2}$) 0.600($6.19\mathrm{e}{-2}$) 0.583   ($6.19\mathrm{e}{-1}$) 0.448   ($7.41\mathrm{e}{-2}$)
@@
UP [17] UnderPressure $0.936^{\dagger}$($1.62\mathrm{e}{-7}$) $0.954^{\dagger}$($9.05\mathrm{e}{-11}$) $0.945^{\dagger}$($7.56\mathrm{e}{-7}$) $0.896^{\dagger}$($7.16\mathrm{e}{-11}$)
@@
Table 8: Foot contact estimation comparison on the MMVP [22] and UnderPressure [17] datasets. Each entry shows the mean metric value and the paired t-test p-value vs. Ours. Bold indicates the best (highest metric); denotes a statistically significant difference from Ours (p < 0.05).
@@ Appendix E Additional Stability Component Estimation Evaluation
Table 9: Comparison of FootFormer (Ours) with baselines for CoP and BoS metrics on PSU-TMM100 [19]. Each entry shows mean $\pm$ std averaged across subjects and thresholds, and the paired t-test p-value vs. Ours. Bold indicates the best (lowest error / highest IoU); denotes a statistically significant difference from Ours (p < 0.05).
@@
CoM Dempster [59] $48.54\pm 33.03^{\dagger}$ (6.72e-06) $44.86\pm 7.53^{\dagger}$ (2.34e-06)
CoMNet [20] $18.80\pm 6.66^{\dagger}$ (4.67e-02) $18.31\pm 7.76^{\dagger}$ (4.28e-02)
@@
Table 10: Comparison of FootFormer (Ours) with baselines for CoM estimation on PSU-TMM100 [19]. Each entry reports mean $\pm$ std or median $\pm$ rSTD error and the paired t-test p-value vs. Ours. Bold indicates the best (lowest error); denotes a statistically significant difference from Ours (p < 0.05).
@@ Appendix F Complete Stability Estimation Evaluation
Figure 9: Comparison of CoM-to-CoP and CoM-to-BoS Absolute Error across all 10 subjects and overall in PSU-TMM100 [19]. We report mean and median errors against CoMNet+PNS [20, 19]. Statistical significance values are presented in Table 11.
@@
CoM-CoP CoMNet+PNS [20] $46.00\pm 22.0^{\dagger}$($2.57\mathrm{e}{-02}$) $29.86\pm 13.2^{\dagger}$($3.63\mathrm{e}{-02}$)
@@
CoM-BoS CoMNet+PNS [20] $34.73\pm 21.7^{\dagger}$($4.50\mathrm{e}{-02}$) $19.79\pm 11.7$($1.48\mathrm{e}{-01}$)
@@
Table 11: Comparison of FootFormer (Ours) with CoMNet+PNS [20] on PSU-TMM100 [19]. Each entry shows mean $\pm$ std or median $\pm$ rSTD error and the paired t-test p-value vs. Ours. Bold indicates the best (lowest error); denotes a statistically significant difference from Ours (p < 0.05). Corresponding per-subject results are in Fig. 9.

### v2

@@ 5.4 Foot Contact
Estimating foot contact (FC), or whether a foot or specific parts of the foot are in contact with the ground plane, is essential for applications in locomotion analysis, rehabilitation, graphics, and animation. We compare FootFormer against FPP-Net [mmvp] and UP [UnderPressure] (Table 1) on their respective datasets to assess effectiveness in contact prediction. Table 3 presents the precision (prec.), recall, F1-score, and IoU evaluation scores.
@@
UP[UnderPressure] UnderPressure $0.936^{\dagger}$ $^{\dagger} 0.9450.945^{\dagger}$ $0.896^{\dagger}$
@@
Table 3: Foot contact estimation results. We train and evaluate on the MMVP [mmvp] and UnderPressure [UnderPressure] datasets and compare against their respective baseline models. Bold indicates the best (highest metric); denotes a statistically significant difference from Ours under paired t-test (p < 0.05).
On the MMVP dataset, FootFormer achieves significantly better precision and equivalent F1, IoU, and recall. Moreover, our model achieves statistically significant improvements over UP across all 4 metrics. Exact p-values are provided in the Supplementary Material. Our results for FPP-Net on the MMVP dataset may be superior to those reported in the original paper [mmvp] as we fully retrain and evaluate with OpenPose [openpose] keypoints instead of their original keypose extraction method [jiang2023rtmpose].
@@ 5.5 Stability Components
Our goal is comprehensive stability analysis. To this end, we evaluate estimates of three stability components, CoP, CoM and BoS, that form the foundation for quantifying human postural stability and balance [jian1993trajectory, lugade2011center, Bruijn_2013]. Figure 3 depicts foot pressure with these stability components and two stability components used in their calculation.
@@
Center of Mass (CoM): CoM represents the weighted average position of body mass and is a crucial factor in a person’s ability to maintain balance. Unlike CoP and BoS, CoM is directly regressed from keypoint sequences. We compare against two baseline methods, CoMNet [JesseRehabJournal], a fully connected network (Table 1), and Dempster’s method [winter2009biomechanics, drillis1964body, dempster1967properties], a classical anthropomorphic method that estimates CoM from weighted sums of segmental centers of mass across the body. For CoM evaluation, Figure 4(a,b) presents both mean and median L2 errors reported in millimeters (mm). FootFormer demonstrates statistically significant improvements over both classical (Dempster’s method [winter2009biomechanics, drillis1964body, dempster1967properties]) and learned (CoMNet [JesseRehabJournal]) baselines.
@@
Figure 4: Performance comparison on  PSU-TMM100 [ScottECCV2020]. (a,b) Mean and median CoM L2 error across subjects. (c) CoP L2 error and (d) BoS IoU across varying pressure thresholds. FootFormer achieves statistically significantly better results across all stability components (see Supplementary Material for exact p-values).
Center of Pressure (CoP): CoP is calculated as the weighted mean of the pressure elements in the XY ground plane. Accurate estimation of CoP is essential for understanding balance, as shifts in CoP can indicate changes in stability or an impending need for postural adjustment [jian1993trajectory]. Base of Support (BoS): BoS represents the area under the feet that supports the body. Estimating BoS is critical for determining the boundaries within which an individual can maintain balance [lugade2011center]. We follow [JesseRehabJournal] and perform foot localization to align the predicted and GT pressure maps with the floor plane using 3D-triangulated keypoints from two viewpoints. Foot position is estimated by foreshortening the pressure map based on dorsiflexion and plantar flexion angles, and rotating it according to ankle orientation ensuring spatial consistency.
@@ 5.6 Stability
Moving beyond simple kinematic estimates, the multiple outputs of FootFormer enable us to directly estimate stability. We calculate two popular measures of postural stability (Figure 3). First, we estimate CoM-CoP defined as $\|CoM-CoP\|_{2}$ or the Euclidean distance from the 2D CoM projected onto the floor plane to CoP [jian1993trajectory]. Typically, the further apart these two points are, the greater the potential for becoming unstable. Second, we measure $\|CoM-BoS_{nearest}\|_{2}$ or the Euclidean distance from the 2D CoM to the boundary of the BoS [lugade2011center]. Intuitively, the CoM-BoS captures the magnitude of instability.
@@
CoM-CoP CoMNet+PNS [JesseRehabJournal] $46.00\pm 22.0^{\dagger}$ $29.86\pm 13.2^{\dagger}$
@@
Table 4: Stability quantification results on PSU-TMM100 [ScottECCV2020]. CoM-CoP and CoM-BoS are reported in mm. Bold indicates the best (lowest error); denotes a statistically significant difference from Ours under paired t-test (p < 0.05).
We compare with Scott et al. [JesseRehabJournal] who use CoMNet to estimate CoM and PNS [ScottECCV2020] to regress foot pressure (Table 1). Table 4 follows [JesseRehabJournal] and reports the mean $\pm$ std and median $\pm$ rSTD for CoM-CoP and CoM-BoS error in mm, where rSTD represents robust standard deviation calculated as the median absolute deviation from the median, multiplied by 1.4826 [rousseeuw1993alternatives]. Error is computed as the absolute distance between predicted and ground-truth positions derived from the insole sensors and MoCap system used in PSU-TMM100. FootFormer achieves statistically significantly improvements in both stability metrics compared to the combined multi-model approach of CoMNet+PNS. We believe this validates the efficacy of learning coupled motion dynamics within a unified structure, as our joint optimization approach outperforms separate models trained independently for each component.
@@ 5.7 Ablation Experiment
Pose Embedding: We compare our learnable GCN against a 1D CNN [mmvp, UnderPressure] and a linear MLP [ScottECCV2020]. Temporal Modeling: We replace the STT with a standard transformer and GRU [mmvp] to assess the efficacy of the spatiotemporal attention mechanism. Contact Conditioning: We evaluate the pressure decoder with and without the contact-based gating to investigate the cross-modal alignment benefits.
@@ A.1 Detailed Architecture Specifications
To adapt the UP [UnderPressure] model to PSU-TMM100, we make simple changes to resize the input and final regression layer to fit PSU-TMM100’s pose input and insole pressure maps, respectively; all other network components are maintained as is. Similarly, we modify FPP-Net’s [mmvp] first layer to fit the joint scheme present in the data. We then adapt the network’s pressure and contact regressor to fit the insole shape and contact regions (like that of FootFormer).
@@ Appendix B Additional Foot Pressure Estimation Results
Figure 6: KLD foot pressure estimation results on PSU-TMM100 [ScottECCV2020] across each subject with mean and median values. We train and evaluate on both 2D and 3D detected keypoints. Lower is better. Statistical significance values are presented in Table 5.
We report the full comparison with Scott et al. [ScottECCV2020], UP [UnderPressure], and FPP-Net [mmvp] on the PSU-TMM100 for each subject in Figure 6. Table 5 reports statistical significance values of paired t-tests comparing FootFormer with the baseline models.
@@
PNS [ScottECCV2020] 2.82 $\pm$ 0.86($1.45\mathrm{e}{-04}$) 2.68 $\pm$ 0.94($9.38\mathrm{e}{-03}$)
FPP-Net [mmvp] 1.40 $\pm$ 0.32   ($2.93\mathrm{e}{-01}$) 1.60 $\pm$ 0.48($1.61\mathrm{e}{-02}$)
UP [UnderPressure] 1.45 $\pm$ 0.35   ($6.06\mathrm{e}{-02}$) 1.50 $\pm$ 0.35($1.59\mathrm{e}{-02}$)
@@
Table 5: Comparison of FootFormer (Ours) with baselines on PSU-TMM100 [ScottECCV2020]. Each entry reports mean $\pm$ std KLD and the paired t-test p-value vs. Ours. Bold indicates the best (lowest KLD); denotes a statistically significant difference from Ours (p < 0.05). Per-subject results are shown in Fig. 6.
@@ 11 Full-Body Stretches 2216
  • Foot pressure maps: Recorded using Tekscan F-Scan 7.0 [tekscan_2020] insole sensors at 50 Hz. Each foot has a high-resolution prexel grid of size $60\times 21$, capturing pressure intensities in kilopascals.
  • 2D and 3D body pose: 2D joints are extracted using OpenPose [openpose] BODY25, while 3D joints are reconstructed via stereo triangulation from the two camera views (like in PSU-TMM100).
@@
Figure 7: Qualitative comparison of predicted foot pressure maps across five OM activities for FootFormer (Ours), PNS [ScottECCV2020], FPP-Net [mmvp], and UP [UnderPressure]. Column headers indicate action; row headers indicate source view or prediction method.
@@ C.1 Additional Ordinary Movements Evaluation
PNS [ScottECCV2020] 3.53 $\pm$ 1.23 ($1.65\mathrm{e}{-04}$) 2.59 $\pm$ 0.94 ($4.35\mathrm{e}{-03}$)
FPP-Net [mmvp] 1.52 $\pm$ 0.37 ($6.89\mathrm{e}{-01}$) 1.54 $\pm$ 0.34 ($9.04\mathrm{e}{-01}$)
UP [UnderPressure] 1.69 $\pm$ 0.34 ($1.28\mathrm{e}{-01}$) 1.48 $\pm$ 0.41 ($6.66\mathrm{e}{-01}$)
@@ Appendix D Additional Foot Contact Estimation Results
Table 8 reports statistical significance values of paired t-tests comparing FootFormer with FPP-Net and UP on the MMVP [mmvp] and UnderPressure [UnderPressure] datasets respectively.
@@
FPP-Net [mmvp] MMVP $0.635^{\dagger}$($1.28\mathrm{e}{-2}$) 0.600($6.19\mathrm{e}{-2}$) 0.583   ($6.19\mathrm{e}{-1}$) 0.448   ($7.41\mathrm{e}{-2}$)
@@
UP [UnderPressure] UnderPressure $0.936^{\dagger}$($1.62\mathrm{e}{-7}$) $0.954^{\dagger}$($9.05\mathrm{e}{-11}$) $0.945^{\dagger}$($7.56\mathrm{e}{-7}$) $0.896^{\dagger}$($7.16\mathrm{e}{-11}$)
@@
Table 8: Foot contact estimation comparison on the MMVP [mmvp] and UnderPressure [UnderPressure] datasets. Each entry shows the mean metric value and the paired t-test p-value vs. Ours. Bold indicates the best (highest metric); denotes a statistically significant difference from Ours (p < 0.05).
@@ Appendix E Additional Stability Component Estimation Evaluation
Table 9: Comparison of FootFormer (Ours) with baselines for CoP and BoS metrics on PSU-TMM100 [ScottECCV2020]. Each entry shows mean $\pm$ std averaged across subjects and thresholds, and the paired t-test p-value vs. Ours. Bold indicates the best (lowest error / highest IoU); denotes a statistically significant difference from Ours (p < 0.05).
@@
CoM Dempster [dempster1967properties] $48.54\pm 33.03^{\dagger}$ (6.72e-06) $44.86\pm 7.53^{\dagger}$ (2.34e-06)
CoMNet [JesseRehabJournal] $18.80\pm 6.66^{\dagger}$ (4.67e-02) $18.31\pm 7.76^{\dagger}$ (4.28e-02)
@@
Table 10: Comparison of FootFormer (Ours) with baselines for CoM estimation on PSU-TMM100 [ScottECCV2020]. Each entry reports mean $\pm$ std or median $\pm$ rSTD error and the paired t-test p-value vs. Ours. Bold indicates the best (lowest error); denotes a statistically significant difference from Ours (p < 0.05).
@@ Appendix F Complete Stability Estimation Evaluation
Figure 9: Comparison of CoM-to-CoP and CoM-to-BoS Absolute Error across all 10 subjects and overall in PSU-TMM100 [ScottECCV2020]. We report mean and median errors against CoMNet+PNS [JesseRehabJournal, ScottECCV2020]. Statistical significance values are presented in Table 11.
@@
CoM-CoP CoMNet+PNS [JesseRehabJournal] $46.00\pm 22.0^{\dagger}$($2.57\mathrm{e}{-02}$) $29.86\pm 13.2^{\dagger}$($3.63\mathrm{e}{-02}$)
@@
CoM-BoS CoMNet+PNS [JesseRehabJournal] $34.73\pm 21.7^{\dagger}$($4.50\mathrm{e}{-02}$) $19.79\pm 11.7$($1.48\mathrm{e}{-01}$)
@@
Table 11: Comparison of FootFormer (Ours) with CoMNet+PNS [JesseRehabJournal] on PSU-TMM100 [ScottECCV2020]. Each entry shows mean $\pm$ std or median $\pm$ rSTD error and the paired t-test p-value vs. Ours. Bold indicates the best (lowest error); denotes a statistically significant difference from Ours (p < 0.05). Corresponding per-subject results are in Fig. 9.